I just finish taking a quiz got 4 wrong out of 10.
1. Solve using the addition principle: -8 + y = 17. My answer was y = 25
2. Solve using the multiplication principle: -12x = 72. My answer was x = -6
3. Solve using the multiplication principle:4/5x = 16. My answer was x = 20
4. Using the formula x = 2a + y, solve for y. My answer was y = -2a+x
Thank you, Jeannette

Answer by Alan3354(69443)

I just finish taking a quiz got 4 wrong out of 10.
1. Solve using the addition principle: -8 + y = 17. My answer was y = 25
25 is correct.
-----------------
2. Solve using the multiplication principle: -12x = 72. My answer was x = -6
-6 is correct.
-------------------
3. Solve using the multiplication principle:4/5x = 16. My answer was x = 20
20 is correct, assuming it's (4/5)x and not 4/(5x)
--------------------------
4. Using the formula x = 2a + y, solve for y. My answer was y = -2a+x
That's correct.